OpenStudy (anonymous):

Solve for x. 2x+3/12 = 2x/8 help please!!!

OpenStudy (dtan5457):

\[\frac{ 2x+3 }{ 12 }=\frac{ 2x }{ 8 }\] does this look right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (dtan5457):

Cross multiply. 2x+3(8)=12(2x) Can you solve for x?

OpenStudy (anonymous):

thanks for the help!!! :)